# Lifelong and Universal Machine Learning Potentials for Chemical Reaction Network Explorations

### v1

@@ 4.3 Reference Data for the Lifelong Machine Learning Potential
To increase the accuracy, training on system-specific data can be a solution. To show that an MLP is in principle able to yield the desired accuracy, we trained an lMLP on the PBE data. Moreover, this large data set allows us to evaluate the performance of continual learning algorithms. These algorithms are necessary for the final goal of rolling explorations, which is the standard setting of an exploratory approach towards chemical reactivity as it may face new reactants, new catalysts, new reductants or oxidants, and so forth, that might not have been well represented by the (initial) training data set. Hence, the lMLP has been designed to efficiently and continuously improve on the fly with new training data getting absorbed (and redundant old data getting forgotten) during the exploration [47] .
@@
The lMLP is typically trained on the difference of the total energy and the sum of reference element energies. These reference element energies differ from the element energies of the previous section in that they are determined before training on some reference structures. In this work, the reference element energies $E_{\mathrm{elem}}^{\mathrm{ref}}$ are calculated from a least squares fit of the total energies of H2, CH4, NH3, and H2222O as a function of their stoichiometries. These reference element energies are applied in Table 1 to show the range and standard deviation of the energies without atomic contributions and atomic force components for the benchmark data.
@@
Table 1: Numbers of structure conformations $N_{\mathrm{conf}}$, as well as ranges and standard deviations of energies $E-E_{\mathrm{elem}}^{\mathrm{ref}}$ and atomic force components $F_{\alpha,n}$ are given for the different data sets.
[…]
$N_{\mathrm{conf}}$ 225 543225543225\,543225 543 225 045225045225\,045225 045 294 916294916294\,916294 916
$\left(E-E_{\mathrm{elem}}^{\mathrm{ref}}\right)^{\mathrm{range}}\,/\,\mathrm{meV\,atom}^{-1}$ 1 578.21578.21\,578.21 578.2 4 076.54076.54\,076.54 076.5 6 159.36159.36\,159.36 159.3
$\left(E-E_{\mathrm{elem}}^{\mathrm{ref}}\right)^{\mathrm{std}}\,/\,\mathrm{meV\,atom}^{-1}$ $61.0$ $131.0$ $151.7$
$F_{\alpha,n}^{\mathrm{range}}\,/\,\mathrm{meV}\,\text{\AA}^{-1}$ 8 71787178\,7178 717 29 9132991329\,91329 913 29 8332983329\,83329 833
$F_{\alpha,n}^{\mathrm{std}}\,/\,\mathrm{meV}\,\text{\AA}^{-1}$ $649$ $448$ $560$
[…]
4.4 Iterative Learning vs. Continual Learning
We first evaluate the performance of the continual learning algorithms to demonstrate that lMLPs can, in principle, be applied in rolling CRN explorations. Only afterwards, we consider lifelong adaptive data selection to reduce the amount of training data considered, and then, in Section 4.6, we investigate whether the lMLP also reaches the target accuracy for CRN energies.
@@
In an lMLP-driven exploration, the lMLP is be pre-trained on an initial data set and then continuously trained on additional, initially unknown data that are flowing in. We constructed a reproducible setting for an exploration process in the following way: We ordered the CRN data by their occurrence during the DFT exploration process and split them into $N_{\mathrm{data}}$ equally sized sets. Hence, the structures in the initial set represent reactions between HCN and H2O, while those of the final sets contain reactions between larger and more complex molecules. For simplicity, we started the performance evaluation by training an lMLP from scratch on the first set and then continuously trained one set after the other, utilizing all training data from each set. This training process was monitored and analyzed as follows.
@@
Figure 4: Test RMSEs of (a) energies $E^{\mathrm{test}}$ and (b) atomic force components $F_{\alpha,n}^{\mathrm{test}}$ after training 1 00010001\,0001 000 epochs on the (extended) data. Trainings were carried out on a sequence of $1$, $2$, $4$, $8$, and $16$ data sets. A constant fraction of the training structures in the respective data set was utilized per epoch (scheme (1)). The total number of epochs $N_{\mathrm{epochs}}$ is higher for training in more sets, resembling the exploitation of previously acquired knowledge. In this figure and in Figures 5, 6, and 7, the underlying training data of each graph coincide at a given value of $n_{\mathrm{epoch}}\,N_{\mathrm{epochs}}^{-1}$. Dots represent RMSEs of individual HDNNP ensemble members, lines show their mean, and shaded areas span their range. The black dashed line represents the mean RMSE of training from scratch.
@@
First, we show how much the continuation of lMLP training can improve the accuracy/cost ratio compared to starting the training after each data addition from scratch again. Therefore, we split the CRN data into $N_{\mathrm{data}}=\{1,2,4,8,16\}$ sets and trained each set for 1 00010001\,0001 000 epochs applying the lADS scheme (1) for the number of structures fitted per epoch (see Section 3.2). Hence, for a larger number of sets, the lMLP is trained in total for more epochs. This scenario resembles the frequently occurring case in which an MLP needs to be improved during a study and the previous MLP can be employed at no additional cost. Therefore, training for the same number of epochs with the same fraction of structures provides a comparison of learning from scratch and continual learning at equal effective cost. The accuracy of the test data after the initial training from scratch for each number of sets is highlighted in Figures 4 (a) and (b). The accuracy of continued training is consistently and significantly better than that of training from scratch for the same underlying training data. This trend highlights the benefit of continual learning compared to conventional iterative learning because it confirms that the previously learned expertise of the model can be exploited. The accuracy appears to converge to a lower limit for training in more sets.
[…]
@@
The simplest learning case is training on all structures from the start, i.e., a stationary batch of data. However, this case is unfeasible in many applications that require either active learning or subsequent tasks with additional training data demands during the applications’ progress. Hence, additional data sets or even a continuous stream of new data need to be learned after the initial training phase. Still, training on all data from the start can provide a reference for the maximally achievable accuracy of the lMLP, so that we can assess the quality of the results obtained with continual learning. Hence, we trained lMLPs on different numbers of data sets ($N_{\mathrm{data}}=\{1,2,4,8,16\}$) to go from this simplest learning case to more and more continual learning. To compare at the same absolute cost, the total number of structure evaluations in the training was the same for all cases. Figures 5 (a) and (b) show that the final test errors increase only slightly with more sets of data. Data addition is visible by spikes in the test RMSE, as the new test data can deviate significantly from previously trained data and can therefore lead to large errors. Recovery of accuracy in a small number of steps demonstrates good integration of the additional data. The peak height reduces with more data since the fraction of new data decreases and the probability increases that the necessary information has already been (partially) trained. Consequently, as expected, incremental learning does not yield the same accuracy as training on all data from the beginning. However, in many practical applications, the large cost reduction of continual learning is more important than the small error increase compared to iterative learning that starts for each data addition from scratch again.
@@ 4.5 Lifelong Adaptive Data Selection
Figure 6: Test RMSEs of (a) energies $E^{\mathrm{test}}$ and (b) atomic force components $F_{\alpha,n}^{\mathrm{test}}$ for training the data in $16$ sets according to scheme (2), with (I) random data selection and disabled stability–plasticity balance in the CoRe optimizer ($p_{\mathrm{frozen}}=0$) and (II) lADS and $p_{\mathrm{frozen}}=0.025$. Each set was trained for 1 00010001\,0001 000 epochs. Hence, both options, (I) and (II), required the same number of structure evaluations.
@@
We now study the effect of lADS and the stability–plasticity balance of the CoRe optimizer on the learning process. The stability–plasticity balance adjusts the plasticity of the model weight parameters; that is, it can freeze important parameters to mitigate forgetting. In Figures 6 (a) and (b), we show how the two approaches change the training in $16$ sets compared to the use of random data selection and disabled stability–plasticity balance ($p_{\mathrm{frozen}}=0$). lADS and the stability–plasticity balance improve the final accuracy of the test energies and forces by $78\%$ and $40\%$, respectively. We note that the number of training structure evaluations is the same (scheme (2)) so that the training costs are almost identical. For random data selection, the characteristic incremental learning convergence pattern vanishes after the initial training sets, since training is not focused on integrating the new data. Instead, the variance of the RMSEs of individual HDNNPs increases. In particular, large-error outliers occur more often, following no pattern. lADS yields a more predictable accuracy with an increase in RMSE after each data addition followed by rapid recovery. Despite this initial error increase, the accuracy is steadily better with lADS than with random data selection and disabled stability–plasticity balance.
@@
lADS and the stability–plasticity balance of the CoRe optimizer can even improve accuracy if all data are trained from the start ($N_{\mathrm{data}}=1$). The difference is less pronounced than in the continual learning case, but replacing random data selection with lADS and increasing $p_{\mathrm{frozen}}$ from 00 to $0.025$ significantly reduces $\mathrm{RMSE}\left(E^{\mathrm{test}}\right)$ from $\left(4.8\pm 0.2\right)$ to $\left(3.6\pm 0.2\right)\,\mathrm{meV\,atom}^{-1}$ and $\mathrm{RMSE}\left(F_{\alpha,n}^{\mathrm{test}}\right)$ from $\left(120\pm 2\right)$ to $\left(108\pm 2\right)\,\mathrm{meV}\,\text{\AA}^{-1}$. The reasons for this effect are that the training is more focused on insufficiently represented training data and that the optimizer can balance the importance of model weight parameters. Consequently, we generally recommend the application of both features.
@@
Figure 7: Total number of reference structure conformations $N_{\mathrm{conf}}$ as a function of the training progress $n_{\mathrm{epoch}}N_{\mathrm{epochs}}^{-1}$ (applying lADS scheme (3)). The colors represent the fractions of test data $N_{\mathrm{test}}$, not employed $N_{\mathrm{train}}^{\mathrm{spare}}$ and employed training data $N_{\mathrm{fit}}$, and training data disregarded due to inconsistencies $N_{\mathrm{incon}}$ and due to redundancy $N_{\mathrm{redun}}$. Data assignment to the classes of being employed or not being employed in fitting can interchange in each epoch, while assignments to test or disregarded data are final.
@@
The best accuracy/cost ratio for training was observed when the number of fitted structures per epoch was adjusted based on a constant fraction of the training structures that were still employed for training (scheme (3)). This scheme is also important for a stable and well-balanced assignment of structures to be redundant in training. Figure 7 shows the evolution of the data assignments for training in $16$ sets, whereby each set was trained for 1 00010001\,0001 000 epochs ($(26.7\pm 0.2)\cdot 10^{6}$ force calculations in total). A steady increase of the number of structures sorted out can be observed. In each of the $10$ individual HDNNP trainings, 134 452±511plus-or-minus134452511134\,452$\pm$ 511134 452 ± 511 or $(66.2\pm 0.3)\%$ of the 202 986202986202\,986202 986 training structures were assigned to be redundant in the end. Only $810\pm 43$ or $(0.40\pm 0.02)\%$ of the training structures were assigned to be inconsistent. Hence, the DFT data appear to be of good quality, and the lMLP is able to represent the majority of these data. Consequently, 67 725±538plus-or-minus6772553867\,725$\pm$ 53867 725 ± 538 or $(33.4\pm 0.3)\%$ structures remain for rehearsal in further continued training. In this way, a significant speed-up can be provided compared to training again on all data for mitigating catastrophic forgetting. The reliability of the lADS approach is confirmed by the small variances in these assignments for the different individually trained HDNNPs of the ensemble.

### v2

@@ 4.2 Reference Data for the Lifelong Machine Learning Potential
The CRN benchmark data set can be utilized to validate that an MLP is in principle able to yield the desired chemical accuracy for CRN explorations. Moreover, this large data set allows us to evaluate the performance of continual learning algorithms. These algorithms are necessary for the final goal of rolling explorations of chemical reactivity. The reason is that the exploration may face new reactants, new catalysts, new reductants or oxidants, and so forth, that might not have been well represented by the (initial) training data set. Hence, the lMLP has been designed to efficiently and continuously improve on the fly, with new training data getting absorbed (and redundant old data getting removed) [58]. As a side remark we note that, without loss of generality for lifelong adaptive data selection, this CRN benchmark dataset only requires the MLP to represent four different chemical elements making it accessible to most MLP descriptors.
@@
MLPs are typically trained on the difference of the total energy and the sum of reference element energies. The latter can be determined before training on some reference structures. In this work, the reference element energies $E_{\mathrm{elem}}^{\mathrm{ref}}$ were calculated from a least squares fit of the total energies of H2, CH4, NH3, and H2O as a function of their stoichiometries. These reference element energies are applied in Table 1 to show the range and standard deviation of the energies without atomic contributions for the benchmark data.
@@
Table 1: Numbers of structure conformations $N_{\mathrm{conf}}$, as well as ranges and standard deviations of energies $E-E_{\mathrm{elem}}^{\mathrm{ref}}$ and atomic force components $F_{\alpha,n}$ for the different data sets.
$\text{PBE}-\text{GFN2}$ PBE
$N_{\mathrm{conf}}$ 225 543225\,543 225 045225\,045
$\left(E-E_{\mathrm{elem}}^{\mathrm{ref}}\right)^{\mathrm{range}}\,/\,\mathrm{meV\,atom}^{-1}$ 1 578.21\,578.2 4 076.54\,076.5
$\left(E-E_{\mathrm{elem}}^{\mathrm{ref}}\right)^{\mathrm{std}}\,/\,\mathrm{meV\,atom}^{-1}$ $61.0$ $131.0$
$F_{\alpha,n}^{\mathrm{range}}\,/\,\mathrm{meV}\,\text{\AA }^{-1}$ 8 7178\,717 29 91329\,913
$F_{\alpha,n}^{\mathrm{std}}\,/\,\mathrm{meV}\,\text{\AA }^{-1}$ $649$ $448$
To further increase the transferability and accuracy of the lMLP, we train the lMLP on the energy difference of PBE DFT energies and GFN2 semi-empirical energies. In this approach, which is inspired by AIQM1 [100], the computational efficiency is still at an affordable level. We note that the semi-empirical method can be replaced by another fast base model. However, the method should not restrict the application range (by contrast to what hampers, for instance, non-reactive force fields), and the difference in the potential energy surface should not be too large, since otherwise the benefit vanishes. For example, a very simple Mie potential with a cutoff radius and parameters trained on DFT data did not improve the results. The advantage of this $\Delta$-learning approach can be seen in Table 1 already because the energy and force ranges are significantly smaller than those of the pure PBE data. Smaller ranges can facilitate the achievement of higher absolute prediction accuracy, since the same relative error leads to a smaller absolute error.
For the PBE data, structures with absolute atomic force components larger than $15\,\mathrm{meV}\,\text{\AA }^{-1}$ were disregarded in lMLP training and performance evaluation, as the accurate prediction of highly unstable structures is not relevant in the application for CRNs. We note that this criterion does not exclude high energy structures per se. For example, the forces of transition states are per definition zero. Since we consider force differences in the $\text{PBE}-\text{GFN2}$ data, we had to adjust the value of the criterion for these data. As our intension is to show that $\Delta$-learning can improve the accuracy, we chose a value that leads to a wider range in terms of structure stability. In this way, representing the data with the same accuracy is more difficult. Therefore, we adjusted the threshold to $4.45\,\mathrm{meV}\,\text{\AA }^{-1}$ for $\text{PBE}-\text{GFN2}$ data, excluding less structures for $\text{PBE}-\text{GFN2}$ data than for PBE data.
4.3 Continual Learning vs. Iterative Learning
This and the following two sections evaluate the performance of lMLPs in different aspects: In this section, we compare the results of the continual learning algorithms to iterative learning. In this way, we can demonstrate the advantage of lMLPs to conventional MLPs in rolling CRN explorations. In the next section, we analyze the accuracy and efficiency gain obtained by lADS compared to random data selection, which are caused by training data focus and reduction. In Section 4.5, we validate that the lMLP can reach the target accuracy for CRN energies.
@@
In an lMLP-driven CRN exploration, the lMLP is pre-trained on an initial data set and then continuously trained on additional, initially unknown data that are flowing in. We constructed a reproducible setting for this exploration process in the following way: We ordered the CRN benchmark data by their occurrence during the DFT exploration process and split them into $N_{\mathrm{data}}$ equally sized sets. Hence, the structures in the initial set represent reactions between HCN and H2O, while those of the final sets contain reactions between larger and more complex molecules. For simplicity, we started the performance evaluation by training an lMLP from scratch on the first set and then continuously trained one set after the other, utilizing all training data from each set. This training process was monitored and analyzed. We note that the continual learning could have been started from a pre-trained lMLP as well.
@@
Figure 4: Test RMSEs of (a) energies $E^{\mathrm{test}}$ and (b) atomic force components $F_{\alpha,n}^{\mathrm{test}}$ after training 1 0001\,000 epochs on the (extended) data. Trainings were carried out on a sequence of 11, 22, 44, 88, and $16$ data sets. The total number of epochs $N_{\mathrm{epochs}}$ is higher for training in more sets, resembling the exploitation of previously acquired knowledge. A constant fraction of $\tfrac{1}{30}$ of all training structures in the respective data set was utilized per epoch. In this way, we can compare results obtained with the same number of structure evaluations, whereby we do not count the evaluations required for training the previous MLP. To avoid instabilities in lADS due to the missing adaption of the number of fitted structures per epoch, we applied here $T_{1}=0.75$, $N_{-}=40$, $N_{++}=400$, and $p_{\mathrm{redun}}^{\mathrm{max}}=0.015$. In this figure and in Figures 5 and 6, $n_{\mathrm{epoch}}\,N_{\mathrm{epochs}}^{-1}$ represents a relative scale for the learning curves on the test data. In this figure, the number of underlying training data coincides for the graphs when dots are plotted at the given value of $n_{\mathrm{epoch}}\,N_{\mathrm{epochs}}^{-1}$ for these graphs. These dots represent RMSEs of individual HDNNP ensemble members, lines show their mean, and shaded areas span their range. The black dashed line represents the mean RMSE of training from scratch.
@@
First, we show how much the continuation of lMLP training can improve the accuracy/cost ratio compared to starting the training after each data addition from scratch again. Therefore, we split the CRN data into $N_{\mathrm{data}}=\{1,2,4,8,16\}$ sets and trained each set for 1 0001\,000 epochs. Hence, for a larger number of sets, the lMLP is trained in total for more epochs. This scenario resembles the frequently occurring case in which an MLP needs to be improved during a study and the previous MLP can be employed at no additional cost. Therefore, training for the same number of epochs with the same constant fraction of structures provides a comparison of learning from scratch and continual learning at equal effective cost. The accuracy of the test data after the initial training from scratch for each number of sets is highlighted in Figures 4 (a) and (b). The accuracy of continued training is consistently and significantly better than that of training from scratch for the same underlying training data. This trend highlights the benefit of continual learning compared to conventional iterative learning because it confirms that the previously learned expertise of the model can be exploited. The accuracy appears to converge to a lower limit for training in more sets.
Figure 5: Test RMSEs of (a) energies $E^{\mathrm{test}}$ and (b) atomic force components $F_{\alpha,n}^{\mathrm{test}}$ for training the data in $1$, 22, 44, 88, and $16$ sets. All these trainings performed the same number of structure evaluations to obtain a fair comparison. Out of this reason, we replaced the adaption of the number of fitted structures per epoch in lADS by a constant number of $N_{\mathrm{fit}}=750$ structures and trained in total for $N_{\mathrm{epochs}}=16\,000$ epochs. We readjusted $p_{\mathrm{redun}}^{\mathrm{max}}$ to $0.0125$ due to the change in $N_{\mathrm{fit}}$. Spikes in the graphs originate from data additions. The dots represent RMSEs of individual HDNNP ensemble members and lines show their mean.
@@
The simplest learning case is training on all structures from the start, i.e., a stationary batch of data. However, this case is unfeasible in many applications that often require active learning and/or subsequent tasks emerge during the application’s progress which have additional training data demands. Hence, additional data sets or even a continuous stream of new data need to be learned after the initial training phase. However, training on all data from the start provides a reference for the maximally achievable accuracy of the lMLP, so that we can assess the quality of the results obtained with continual learning. Hence, we trained lMLPs on different numbers of data sets ($N_{\mathrm{data}}=\{1,2,4,8,16\}$) to go from this simplest learning case to more and more continual learning. To compare at the same absolute cost, the total number of structure evaluations in the training was the same for all cases. Figures 5 (a) and (b) show that the final test errors increase only slightly with more sets of data. Data addition is visible by spikes in the test RMSE, as the new test data can deviate significantly from previously trained data and can therefore lead to large errors. Recovery of accuracy in a small number of steps demonstrates good integration of the additional data. The peak height reduces with more data since the fraction of new data decreases and the probability increases that the necessary information has already been (partially) trained. In conclusion, incremental or continual learning does not yield the same accuracy as training on all data from the beginning as expected. However, in many practical applications, the large cost reduction of continual learning is more important than the small error increase compared to iterative learning that starts for each data addition from scratch again.
@@ 4.4 Lifelong Adaptive Data Selection
Figure 6: Test RMSEs of (a) energies $E^{\mathrm{test}}$ and (b) atomic force components $F_{\alpha,n}^{\mathrm{test}}$ for training the data in $16$ sets, with (I) random data selection and disabled stability–plasticity balance in the CoRe optimizer ($p_{\mathrm{frozen}}=0$) and (II) lADS and $p_{\mathrm{frozen}}=0.025$. Analogous to Figure 5, a constant number of fitted structures per epoch (including the readjustment of $p_{\mathrm{redun}}^{\mathrm{max}}$) was employed for a fair comparison based on the same number of structure evaluations. Each of the $16$ sets was trained for 1 0001\,000 epochs. The dots represent RMSEs of individual HDNNP ensemble members and lines show their mean.
@@
In this section, we study (I) the effect of continual learning approaches (lADS and the stability–plasticity balance of the CoRe optimizer) on the learning process and (II) the reduction of training data obtained by lADS. The stability–plasticity balance of the CoRe optimizer can freeze important model weight parameters to mitigate forgetting. In Figures 6 (a) and (b), we show how lADS and the stability–plasticity balance change the learning curve on test data for training in $16$ sets compared to the use of random data selection and disabled stability–plasticity balance ($p_{\mathrm{frozen}}=0$). The two continual learning approaches improve the final accuracy of the test energies and forces by $78\%$ and $40\%$, respectively. We note that the number of training structure evaluations is the same so that the training costs are almost identical. For random data selection, the characteristic incremental learning convergence pattern vanishes after the initial training sets, since training is not focused on integrating the new data. Instead, the variance of the RMSEs of individual HDNNPs increases. In particular, large-error outliers occur more often, following no pattern. lADS yields a more predictable accuracy with an increase in RMSE after each data addition followed by rapid recovery. Despite this initial error increase, the accuracy is steadily better with lADS than with random data selection and disabled stability–plasticity balance.
@@
lADS and the stability–plasticity balance can even improve the accuracy if all data are trained from the start ($N_{\mathrm{data}}=1$). The difference is less pronounced than in the continual learning case, but replacing random data selection with lADS and increasing $p_{\mathrm{frozen}}$ from 0 to $0.025$ significantly reduces $\mathrm{RMSE}\left(E^{\mathrm{test}}\right)$ from $\left(4.8\pm 0.2\right)$ to $\left(3.6\pm 0.2\right)\,\mathrm{meV\,atom}^{-1}$ and $\mathrm{RMSE}\left(F_{\alpha,n}^{\mathrm{test}}\right)$ from $\left(120\pm 2\right)$ to $\left(108\pm 2\right)\,\mathrm{meV}\,\text{\AA }^{-1}$. The reasons for this effect are that the training is more focused on insufficiently represented training data and that the optimizer can balance the importance of model weight parameters. Consequently, we generally recommend the application of both features.
@@
Figure 7: Total number of reference structure conformations $N_{\mathrm{conf}}$ as a function of the training progress $n_{\mathrm{epoch}}N_{\mathrm{epochs}}^{-1}$. The colors represent the fractions of test data $N_{\mathrm{test}}$, not employed $N_{\mathrm{train}}^{\mathrm{spare}}$ and employed training data $N_{\mathrm{fit}}$, and training data disregarded due to inconsistencies $N_{\mathrm{incon}}$ and due to redundancy $N_{\mathrm{redun}}$. Data assignment to the classes of being employed or not being employed in fitting can interchange in each epoch, while assignments to test or disregarded data are final. The number of fitted structures per epoch was adjusted to the number of training structures that were not sorted out (as in Algorithm 1).
@@
The best accuracy/cost ratio for training was observed when the number of fitted structures per epoch was adjusted based on a constant fraction of the training structures that were still employed for training. This adjustment is also important for a stable and well-balanced assignment of structures to be redundant in training. Figure 7 shows the evolution of the data assignments for training in $16$ sets, whereby each set was trained for 1 0001\,000 epochs ($(26.7\pm 0.2)\cdot 10^{6}$ force calculations in total). A steady increase of the number of structures sorted out can be observed. In each of the $10$ individual HDNNP trainings, $134\,452\pm 511$ or $(66.2\pm 0.3)\%$ of the 202 986202\,986 training structures were in the end classified as redundant. Only $810\pm 43$ or $(0.40\pm 0.02)\%$ of the training structures were classified as inconsistent. Hence, the DFT data appear to be of good quality, and the lMLP is able to represent the majority of these data. Consequently, $67\,725\pm 538$ or $(33.4\pm 0.3)\%$ of the structures remain for rehearsal in further continued training. In this way, a significant speed-up can be provided compared to training again on all data for mitigating catastrophic forgetting. The reliability of the lADS approach is confirmed by the small variances in these assignments for the different individually trained HDNNPs of the ensemble.